Circular No. 8876

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 S2 (LEMMON) An apparently asteroidal object discovered on CCD images taken with

the 1.5-m reflector in the course of the Mt. Lemmon survey (discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to show cometary appearance by two observers: Q.-z. Ye (Department of Atmosphere Science, Sun Yat-sen University, Guangzhou, China) writes that three stacked 60-s CCD exposures taken with the Lulin Observatory 0.41-m Ritchey-Chretien reflector on Sept. 27.85 UT show a very obvious coma of diameter ∼ 5′′.5 with a short fan-like tail 6′′ long towards p.a. ∼ 180^{o}–270^{o}. J. W. Young reports that his images taken with the Table Mountain 0.61-m f/16 Cassegrain reflector on Oct. 3.45–3.53 show a 6′′-diameter, very round, bright coma with no hint of a tail.

2007 UT  α_{2000}  δ_{2000}  Mag.
Sept.25.49148  8^{h}21^{m}53.^{s}89  +19^{o}54′08′′.5  18.9

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-T11. This comet’s orbit is probably of intermediate or short period.

T = 2007 Oct. 12.007 TT  ω = 179.^{o}725
Ω = 296.835 }2000.0
q = 6.21749 AU  i = 15.669

V455 ANDROMEDAE E. Broens, Mol, Belgium, reports following visual magnitude estimates

of V455 And (cf. IAUC 8868) on behalf of the Werkgroep Veranderlijke Sterren, Vereniging Voor Sterrenkunde: Sept. 11.811 UT, 10.7 (D. Vansteelant, Varsenare, Belgium); 12.860, 10.6 (H. Hautecler, Boutersem, Belgium); 13.868, 10.6 (A. Baillien, Tongeren, Belgium), 14.863, 10.8 (Vansteelant); 15.808, 11.2 (A. Diepvens, Balen, Belgium); 15.894, 10.8 (Hautecler); 16.806, 11.0 (Baillien); 18.860, 11.3 (P. Cloesen, Beringen, Belgium); 21.891, 11.4 (Vansteelant); 22.791, 12.0 (Diepvens); 23.857, 12.9 (Cloesen); 24.816, 13.2 (E. Muyllaert, Oostende, Belgium); 25.802, 13.1 (Vansteelant). CCD V magnitudes by M. Martignoni, Magnago, Italy: Sept. 11.972, 10.22; 12.933, 10.30; 15.035, 10.58; 15.977, 11.01; 19.052, 11.32; 20.023, 11.48; 23.990, 12.66.

2007 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green